FIND THE RANGE OF VALUES OF a(an element of R)for which z+a*(mod of (z-1))+2i=0,where z=x+iy has a solution.also find the solution
Answer by venugopalramana(3286)

FIND THE RANGE OF VALUES OF a(an element of R)for which z+a*(mod of (z-1))+2i=0,where z=x+iy has a solution.also find the solution
I DONT THINK THE PROBLEM IS PROPER.
FOR A COMPLEX NUMBER TO BE ZERO REAL AND IMAGINARY PARTS ARE TO BE ZERO
Z=X+IY
MOD OF Z-1 IS REAL ,A IS REAL..HENCE A*|Z-1| IS REAL ..
2I IS IMAGINARY
HENCE
IMAGINARY PART = IY+2I=0....Y=-2...HENCE THIS CANT BE TRUE FOR ANY VALUE OF Y OR INTURN FOR ANY VALUE OF Z.
REAL PART =X+A*SQRT[(X-1)^2+Y^2]=0...PUTTING Y=-2
A=-X/SQRT.[(X-1)^2+4]
THEN A DEPENDS ON VALUE OF X